OpenStudy (anonymous):

A bacteria culture begins with 15 bacteria which double in amount at the end of every hour. How many bacteria exist in the culture at the end of the 12th hour?

OpenStudy (anonymous):

@SolomonZelman @kc_kennylau

OpenStudy (anonymous):

@Disco619

OpenStudy (diyadiya):

After every hour it doubles 1st hour- 15*2 =30 2nd hour-(15*2)*2=15*\(15*2^{2}\)=60 3rd hour-(15*2*2)*2=\(15*2^{3}\)=120 4th hour-(15*2*2*2)*2=\(15*2^{4}\)=240 so after 12th hour \(15*2^{12}\)

OpenStudy (disco619):

Im not really good at these "arithmetic progression" equations but the answer is: 61440

OpenStudy (anonymous):

@kc_kennylau where does the n-1 come from?

OpenStudy (diyadiya):

@kc_kennylau after first hour it will double

OpenStudy (disco619):

We are not gonna count the 15 bacteria. They don't come in the 12 hours. They are the beginning

OpenStudy (dan815):